OpenStudy (anonymous):

What is the quotient: (65y3 + 15y2 – 25y) ÷ 5y

OpenStudy (anonymous):

\[\frac{65y^3 + 15y^2 – 25y}{5y}=\frac{65y^2}{5y}+\frac{15y^2}{5y}+\frac{-25y}{5y}\] that should be a bit easier to find.

OpenStudy (anonymous):

so i think he answer is 13 y^2 - 3y+5

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

oh alright thnx for the help

OpenStudy (shane_b):

Wait...I was looking at @tomo's response and he had a typo in his conversion. Anyway, the final answer should have been 13y^2+3y-5

OpenStudy (anonymous):

ill be looking for my medal soon lol

OpenStudy (anonymous):

it should have been 65y^3 instead of 65y^2 on the first term

OpenStudy (shane_b):

Yes

OpenStudy (shane_b):

Anyway, I believe @chase3 gets the idea

OpenStudy (anonymous):

yea i got it